# Composable Constraint Models for Permutation Enumeration

### v3

@@ 1 Introduction
We believe our library of models, which can be easily composed in a "pick and mix" fashion, will further the research in permutation patterns by allowing for more efficient computational experimentation and exhaustive search as the constraint programming approach avoids the need for the traditional "generate-and-test" approach, which generates intermediate assignments that only satisfy some of the properties for later filtering.
This "generate-and-test" approach, which requires specialised algorithms to be created to find all permutations which satisfy some given property, can work very well for that property but this approach is difficult to extend and requires considerable expertise in programming to add new properties. On the other hand, combining CP models can immediately allow highly efficient searching for a combination of properties, and our experience is adding new patterns in CP is significantly easier than create new bespoke state-of-the-art programs for searching for new patterns.
@@ 2.1.1 Constraint Types
Iterating over elements (Line 12):
@@
Matrix constraints (Line 13):
@@
Conditional constraints (Line 14):
@@
Pattern avoidance (Line 15):
@@
Fixed values (Line 16):
@@
Listing 2: Essence code which represents classic containment
@@
1\$ A single, random permutation of length 4, containing 21
@@
5\$ 10 random permutations of length 4, containing 21
6[{"perm": [1, 2, 4, 3]}, {"perm": [1, 3, 2, 4]}, {"perm": [1, 3, 4, 2]}, {"perm": [1, 4, 2, 3]}, {"perm": [1, 4, 3, 2]},
7 {"perm": [2, 1, 3, 4]}, {"perm": [2, 1, 4, 3]}, {"perm": [2, 3, 1, 4]}, {"perm": [2, 3, 4, 1]}, {"perm": [2, 4, 1, 3]}]
@@ 3.1 Pattern types
Listing 4: Essence code which represents classic containment
@@
9\$ We use a matrix to represent the permutation in two dimensions, this is not really needed for classic patterns, but to allow for the code of the different
10\$ patterns to be composable we have added this here as well.
@@
14 (forAll n1, n2 : int(1..|pattern|) , n1 < n2 . pattern(n1) < pattern(n2) <-> perm(ix[n1]) < perm(ix[n2])))
@@
Listing 5: Essence code which represents vincular containment
@@
12 (forAll n1, n2 : int(1..|pattern|) , n1 < n2 . (pattern(n1) < pattern(n2) <-> perm(ix[n1]) < perm(ix[n2]))) .
13\$ Then we look in the matrix representation if the adjacency for this pattern is not violated.
@@
Listing 6: Essence code which represents bivincular containment
@@
6\$ Creating a padded version of perm, where position 0 contains the value 0 and position length+1 contains the value length+1, to be able to
7\$ follow the convention.
@@
31 (forAll bar in val_bars . permPadded[ix[patterninv[bar]]]+1 = permPadded[ix[patterninv[bar+1]]]))
  • $\sigma(i_{1})\ldots\sigma(i_{k})$ is order isomorphic to $\pi$
Listing 7: Essence code which represents mesh containment
@@
6\$ Creating a padded version of perm, where position 0 contains the value 0 and position length+1 contains the value length+1, to be able to
7\$ follow the convention.
@@
Listing 8: Essence code which represents classic avoidance
@@
9\$ We use a matrix to represent the permutation in two dimensions, this is not really needed for classic patterns, but to allow for the code of the different
10\$ patterns to be composable we have added this here as well. While the classic pattern constraint stays the same we are now negating the whole thing
11\$ (see the ! before the exists statement). So we are looking for a permutation in which it is not possible to find the pattern in.
@@
13\$ We now look constraint perm to contain an order isomorphic copy of pattern.
@@
15 (forAll n1, n2 : int(1..|pattern|) , n1 < n2 . pattern(n1) < pattern(n2) <-> perm(ix[n1]) < perm(ix[n2])))
@@
Listing 9: Essence code which represents mesh avoidance
@@
6\$ Creating a padded version of perm, where position 0 contains the value 0 and position length+1 contains the value length+1, to be able to
7\$ follow the convention.
@@
19\$ Look for all places where the pattern can occur, in each we need to violate the mesh. This changes from an ’exists’ to a ’for all’
@@
25\$ If we find the pattern, make sure there is at least one value in some cell of the mesh
@@ 3.2 Permutation Properties
Listing 10: Essence code which represents simple permutations
@@
4\$ perm is a simple permutation, subperm is a set of contiguous indices.
5such that
@@
7\$ There is a split where there is an interval on the "left" where the maximal point is below the lowest point in the interval on the "right".
@@
7\$ There is a split where there is an interval on the "left" where the minimal point is above the highest point in the interval on the "right".
@@
Listing 14: Essence code which represents derangements
@@
5\$ None of the indices are fixed points.
6 forAll i : int(1..length) . perm(i) != i
@@
Listing 15: Essence code which represents nonderangements
@@
6 exists i : int(1..length) . perm(i) = i
@@
Listing 16: Essence code which represents involutions
@@
6\$ perm * perm = identity permutation (increasing permutation)
7 forAll i : int(1..length) . perm(perm(i)) = i
@@
Listing 17: Essence code which represents permutations with parity
@@
5\$ Check for odd/even by doing modulo 2 arithmetic.
6 forAll i : int(1..length) . (perm(i) % 2) = i % 2
@@ 3.3 Permutation Statistics
Listing 18: Essence code which represents all statistics
@@
4find inversionCount : int(0..length**2) \$ We need to each statistic it a reasonable upper bound
5such that
6\$ For each occurence of a larger value is found before a smaller one, we add a 1 to a list and we then sum them all up.
7 inversionCount = sum([ 1 | i,j : int(1..length), i < j, perm(i) > perm(j)])
8
9such that
10\$ For each occurence of a descent, we add a 1 to a list and we then sum them all up.
11 descentCount = sum([ 1 | i : int(1..length-1), perm(i) > perm(i+1)])
12
13find ascentCount : int(0..length)
14such that
15\$ For each occurence of an ascent, we add a 1 to a list and we then sum them all up.
16 ascentCount = sum([ 1 | i : int(1..length-1), perm(i) < perm(i+1)])
@@
18find excedanceCount : int(0..length)
19such that
20\$ For each occurence of an excedance, we add a 1 to a list and we then sum them all up.
21 excedanceCount = sum([ 1 | i : int(1..length), perm(i) > i])
22
23find majorIndex : int(0..length**2)
24such that
25\$ We follow the sum by the definition of Major Index.
26 majorIndex = sum([ i | i : int(1..length-1), perm(i) > perm(i+1) ])
@@
7 nInversions = sum([ 1 | i,j : int(1..length), i < j, perm(i) > perm(j)])

### v4

@@ Abstract
keywords:
Constraint Modelling, Permutation Pattern, Enumeration
@@ 1 Introduction
We believe our library of models, which can be easily composed in a ”pick and mix” fashion, will further the research in permutation patterns by allowing for more efficient computational experimentation and exhaustive search as the constraint programming approach avoids the need for the traditional ”generate-and-test” approach, which generates intermediate assignments that only satisfy some of the properties for later filtering.
This ”generate-and-test” approach, which requires specialised algorithms to be created to find all permutations which satisfy some given property, can work very well for that property but this approach is difficult to extend and requires considerable expertise in programming to add new properties. On the other hand, combining CP models can immediately allow highly efficient searching for a combination of properties, and our experience is adding new patterns in CP is significantly easier than create new bespoke state-of-the-art programs for searching for new patterns.
@@ 2.1.1 Constraint Types
Iterating over elements (Line 10):
@@
Matrix constraints (Line 11):
@@
Conditional constraints (Line 12):
@@
Pattern avoidance (Line 13):
@@
Fixed values (Line 14):
@@
Listing 2: Essence code which represents classic containment
@@
1# A single, random permutation of length 4, containing 21
@@
5# 10 random permutations of length 4, containing 21
6[{"perm": [1, 2, 4, 3]}, {"perm": [1, 3, 2, 4]}, {"perm": [1, 3, 4, 2]}, {"perm": [1, 4, 2, 3]},
7 {"perm": [1, 4, 3, 2]}, {"perm": [2, 1, 3, 4]}, {"perm": [2, 1, 4, 3]}, {"perm": [2, 3, 1, 4]},
8 {"perm": [2, 3, 4, 1]}, {"perm": [2, 4, 1, 3]}]
@@ 2.3 Constraint Solving
Once the model has been translated into a solver-compatible format, the solver takes over to search for solutions using a combinatorial search algorithm, most commonly a form of backtrack search. This process systematically assigns values to variables while checking that all constraints are satisfied.
@@ 3.1 Pattern types
9\$ We use a matrix to represent the permutation in two dimensions, this is not really needed for classic patterns,
10\$ but to allow for the code of the different patterns to be composable we have added this here as well.
@@
14 (forAll n1, n2 : int(1..|pattern|) , n1 < n2 .
15 pattern(n1) < pattern(n2) <-> perm(ix[n1]) < perm(ix[n2])))
Listing 4: Essence code which represents classic containment
@@
12 (forAll n1, n2 : int(1..|pattern|) , n1 < n2 .
13 (pattern(n1) < pattern(n2) <-> perm(ix[n1]) < perm(ix[n2]))) .
14\$ Then we look in the matrix representation if the adjacency for this pattern is not violated.
@@
Listing 5: Essence code which represents vincular containment
@@
6\$ Creating a padded version of perm, where position 0 contains the value 0 and position length+1 contains the value length+1,
7\$ to be able to follow the convention.
@@
31 (forAll bar in val_bars . permPadded[ix[patterninv[bar]]]+1 =
32 permPadded[ix[patterninv[bar+1]]]))
Listing 6: Essence code which represents bivincular containment
@@
6\$ Creating a padded version of perm, where position 0 contains the value 0 and position length+1 contains the value length+1,
7\$ to be able to follow the convention.
@@
Listing 7: Essence code which represents mesh containment
  • $\sigma(i_{1})\ldots\sigma(i_{k})$ is order isomorphic to $\pi$
@@
9\$ We use a matrix to represent the permutation in two dimensions, this is not really needed for classic patterns,
10\$ but to allow for the code of the different patterns to be composable we have added this here as well.
11\$ While the classic pattern constraint stays the same we are now negating the whole thing (see the ! before the exists statement)
12\$ So we are looking for a permutation in which it is not possible to find the pattern in.
@@
14\$ We now look constraint perm to contain an order isomorphic copy of pattern.
@@
16 (forAll n1, n2 : int(1..|pattern|) , n1 < n2 .
17 pattern(n1) < pattern(n2) <-> perm(ix[n1]) < perm(ix[n2])))
Listing 8: Essence code which represents classic avoidance
@@
6\$ Creating a padded version of perm, where position 0 contains the value 0 and position length+1 contains the value length+1,
7\$ to be able to follow the convention.
@@
19\$ Look for all places where the pattern can occur, in each we need to violate the mesh.
20\$ This changes from an ’exists’ to a ’for all’
@@
26\$ If we find the pattern, make sure there is at least one value in some cell of the mesh
@@
Listing 9: Essence code which represents mesh avoidance
@@ 3.2 Permutation Properties
4\$ Simple permutations do not contain intervals.
5\$ An interval is a set of contiguous indices such that the values are also contiguous.
6\$ perm is a simple permutation, subperm is a set of contiguous indices.
7such that
@@
Listing 10: Essence code which represents simple permutations
@@
7\$ There is a split where there is an interval on the ”left” where the maximal point is below the lowest point in the interval on the ”right”.
@@
7\$ There is a split where there is an interval on the ”left” where the minimal point is above the highest point in the interval on the ”right”.
@@
5 forAll i : int(1..length) .
6\$ None of the indices are fixed points.
7 perm(i) != i
Listing 14: Essence code which represents derangements
@@
6 exists i : int(1..length) .
7 perm(i) = i
Listing 15: Essence code which represents nonderangements
@@
6 forAll i : int(1..length) .
7\$ perm * perm = identity permutation (increasing permutation)
8 perm(perm(i)) = i
Listing 16: Essence code which represents involutions
@@
5 forAll i : int(1..length) .
6\$ Check for odd/even by doing modulo 2 arithmetic.
7 (perm(i) % 2) = i % 2
Listing 17: Essence code which represents permutations with parity
@@ 3.3 Permutation Statistics
4\$ the inversion count
5find inversionCount : int(0..length**2) \$ We need to each statistic it a reasonable upper bound
6such that
7 inversionCount =
8\$ For each occurence of a larger value is found before a smaller one, we add a 1 to a list and we then sum them all up.
9 sum([ 1 | i,j : int(1..length), i < j, perm(i) > perm(j)])
10
11\$ the descent count
12find descentCount : int(0..length)
13such that
14 descentCount =
15\$ For each occurence of a descent, we add a 1 to a list and we then sum them all up.
16 sum([ 1 | i : int(1..length-1), perm(i) > perm(i+1)])
@@
18\$ the ascent count
19find ascentCount : int(0..length)
20such that
21 ascentCount =
22\$ For each occurence of an ascent, we add a 1 to a list and we then sum them all up.
23 sum([ 1 | i : int(1..length-1), perm(i) < perm(i+1)])
24
25\$ the excedance count
26find excedanceCount : int(0..length)
27such that
28 excedanceCount =
29\$ For each occurence of an excedance, we add a 1 to a list and we then sum them all up.
30 sum([ 1 | i : int(1..length), perm(i) > i])
31
32\$ the majorIndex
33find majorIndex : int(0..length**2)
34such that
35 majorIndex =
36\$ We follow the sum by the definition of Major Index.
37 sum([ i | i : int(1..length-1), perm(i) > perm(i+1) ])
Listing 18: Essence code which represents all statistics
@@
7 nInversions =
8 sum([ 1 | i,j : int(1..length), i < j, perm(i) > perm(j)])